# VizAbility: Enhancing Chart Accessibility with LLM-based Conversational Interaction

### v3

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Visualization systems and tools
@@ 1. Introduction
Numerous assistive technologies have been developed to allow BLV users to access visualizations using sensory modalities other than vision (Kim et al., 2021). Tactile visualizations can provide a tangible representation of data while necessitating specialized hardware such as haptic displays  (Paneels and Roberts, 2009) and embossing machines (Engel and Weber, 2017a). On the other hand, sonification can enable users to discern trends and anomalies through sound (Wang et al., 2022), but it is typically limited to single-series data. Traditional methods for adapting web visualizations for screen readers include data tables and alternative text (Kim et al., 2021), while these methods often diminish the inherent advantages of data visualizations. Keyboard-based chart navigation (Weninger et al., 2015; Elavsky et al., 2024; Thompson et al., 2023; Zong et al., 2022) have emerged as an alternative solution. However, orienting oneself and navigating within complex chart encoding structures pose challenges for efficient data exploration (Kim et al., 2023a).
This work introduces VizAbility, a novel approach to augment keyboard navigation of chart content with conversational interaction (Figure 1). First, we use Olli’s tree view (referenced in (Blanco et al., 2022)) to create a keyboard-navigable text representation of a chart. Next, we enhance this tree view with an LLM-based question-and-answer module for addressing on-demand queries. These natural language queries allow users to understand the chart without needing to navigate and mentally synthesize different parts of the chart to derive insights. VizAbility uses the user’s position within the tree view to efficiently respond to both visual and analytical queries, facilitating the exploration of data trends and visual patterns. Moreover, VizAbility can handle contextual queries, providing background information about the chart, tailored specifically for the needs of BLV individuals (Kim et al., 2023b). We based these query types on actual questions asked by BLV people in previous studies (Kim et al., 2020, 2023b). Additionally, it can manage navigation queries, allowing users to control their position through verbal commands.
@@
We carried out quantitative assessments to evaluate the question & answering pipeline. We evaluated response quality using a combined dataset of 979 real BLV user questions derived from previous research (Kim et al., 2023b) and 48 synthetically generated navigation queries. Splitting the dataset, 80% was used for testing and 20% for validation. Our query classification achieved an accuracy of 87.39%. The final response evaluation involved a manual assessment by two researchers, followed by a more scalable LLM-based evaluation using GPT4. Both the human and GPT4 assessments followed a 5-point Likert Scale that assigned a value ranging from “Very Poor” to “Very Good” depending on the response’s coherence to the ground truth. For the human evaluation, $77.36\%$ and $5.14\%$ of the responses were rated as “Very Good’ and “Good” respectively, contributing to an overall rate of $82.50\%$. We computed Kendall’s $\tau$ score ($=0.5526$, $p<0.001$) to assess the consistency between the human and GPT4 assessment methods, observing significant alignment. As a baseline comparison, we used the GPT4-based evaluation to assess responses generated by GPT-4 with vision (GPT-4V) on the same test data. We observed the performance of GPT-4V was poorer, with only 27.96% and 10.10% of responses being designated as “Very Good” and “Good” respectively.
@@
  • Design and development of VizAbility, which incorporates an LLM-based question-and-answer module to enhance keyboard navigation of chart content for screen reader users
@@
The VizAbility source code and dataset are available at [redacted for anonymous review].
@@ 2.1. Accessibility Systems for Data Visualization
Screen readers, utilizing text/speech modalities, stand as the predominant assistive technology, particularly for navigating web content. The go-to accessibility techniques for screen readers encompass alternative text and data tables. Yet, these strategies often reduce data visualizations to brief descriptions or mere numbers, undermining their inherent advantages. An alternative approach involves crafting keyboard-navigable text descriptions derived from the chart’s structure. A select group of data visualization tools and toolkits, such as HighCharts (Highcharts, 2022) and amCharts (amc, [n. d.]), offer some degree of this navigation and customization (Kim et al., 2023a). In recent times, several systems have advanced their navigation capabilities, representing charts as traversable graph structures (Weninger et al., 2015; Godfrey et al., 2018; Zong et al., 2022; Thompson et al., 2023; Elavsky et al., 2024).
@@
To address the limitations of relying on a single sensory modality, multi-sensory perception is frequently utilized. A prevalent strategy involves merging verbal (speech) cues with non-verbal ones, such as sonification, tactile graphics, and haptic feedback. Examples include offering on-demand audio descriptions of touched elements (Goncu and Marriott, 2011; Gardner and Bulatov, [n. d.]; Landau and Gourgey, 2001) or pairing sonification with speech or screen readers (Thompson et al., 2023; Taibbi et al., 2014). However, these solutions often necessitate specialized software and hardware, especially for interactive tactile support, making them expensive to implement.
@@ 2.2. Question & Answering Systems for Data Visualization
Recent studies have begun to focus on data visualization images (Hoque et al., 2022). For example, FigureQA (Kahou et al., 2017) offers a corpus tailored for yes/no questions, such as “Is Light Gold less than Periwinkle?”. Conversely, DVQA (Kafle et al., 2018) expands its purview to encompass questions about chart structure (“are the bars horizontal?”), data retrieval (“what percent of people prefer A?”), and reasoning (“Is A preferred more than B?”). While both FigureQA and DVQA rely on synthetically generated charts, PlotQA introduces a large-scale dataset of real-world scientific plots. Unlike the templated questions of the aforementioned datasets, ChartQA delivers human-composed questions, enhanced using LLMs (Masry et al., 2022). These models predominantly process pixel images as input. For instance, they extract data tables and other image features (Masry et al., 2022; Kantharaj et al., 2022), feeding them into vision and language task models (Cho et al., 2021). Consequently, their accuracy largely hinges on their image processing capabilities, often leading to suboptimal results (e.g., failing to recover data values due to the absence of data labels). In a different approach, Kim et al.(Kim et al., 2020) proposed a system that not only answers questions but also provides explanations, operating on Vega-lite(Satyanarayan et al., 2017) instead of images. All the current question-answering systems are limited to basic visualization types like bar, line, and pie charts.
@@ 3. VizAbility Design Decisions
Bridging the perceptual gap between BLV and sighted individuals requires a deep understanding of what the chart looks like. While some blind individuals may not prioritize visual encoding information (Thompson et al., 2023; Lundgard and Satyanarayan, 2022), previous research indicates that navigating charts based on their visual encoding helps BLV users gain a clearer visual understanding (Kim et al., 2023a). In this work, we use Olli (Blanco et al., 2022) to generate a keyboard-navigable text representation of charts.
@@
D7: Follow POUR accessibility principles
The W3C Web Accessibility Initiative specifies four essential principles for web accessibility: Perceivable, Operable, Understandable, and Robust (wai, 2023). Ensuring these basic principles while developing a new assistive technology can be easily overlooked but is important to enhance its overall utility. We strive to adhere to these principles in the design of VizAbility. Examples include reaffirming user queries (perceivable); suggesting cold-start queries (operable), indicating delays in responses (understandable), and allowing both speech and text inputs to formulate queries (robust).
@@ 4.4. Rapid Chart Probing via Conversational Interaction
Keyboard navigation helps blind users grasp the chart’s visual structure and data, but can be cumbersome for higher-level comprehension, such as computing aggregates or understanding overall visuals. We integrate LLMs for intuitive chart question-answering and enhance interaction by combining keyboard navigation with speech inputs.
@@ 4.4.1. Data Set
The ground truths for the testing and validation sets were manually generated by two researchers independently and then merged by resolving conflicts. The process involved reading charts, calculating numbers, and searching for information online. Throughout the generation process, we emphasized verboseness. For instance, the ground truth response to the question “What is the vaccination rate of South Africa” is “The vaccination rate for South Africa is 36%”, as opposed to the more concise “36%”. To prioritize the model’s ability to translate ambiguous user wording into precise nodes within the treeview, the ground truth responses for navigation queries consist solely of starting and ending nodes.
@@ 4.4.2. Supported Query Types
Analytical queriesD2 primarily focus on understanding the underlying data, such as “Is Africa the country that needs the vaccine the most?” or “What is the highest positive anomaly?” Visual queriesD2 relate to visual encoding information or demand visual interpretation, exemplified by questions like “What color is North America?” or “Is the line fluctuating?” Analytical and visual queries are not entirely distinct; visual queries often necessitate data interpretation, as in “Which country exhibits the darkest shades for both the lowest and highest values?”. However, we continue to differentiate these query types to communicate to LLMs that a query involves information encoded visually, which is specific to the problem of chart and question answering.
@@ 4.4.4. Query-Specific Prompting
D2
Figure 6 illustrates our approach to handling analytical and visual queries. To circumvent the predefined token limit of the LLM, we consolidate the transformed data extracted from the Vega View (veg, [n. d.]) into an external CSV file. This file is then processed by LangChain’s CSV Agent (csv, [n. d.]), which operates in the background. Under the hood, this agent leverages the Pandas DataFrame agent, subsequently executing Python code generated by the LLM. We purposefully avoid including the entire raw dataset, recognizing that it might differ from the final view data. Often, the agent can get stuck in an infinite loop of thinking. To prevent this issue, we have implemented a time constraint. If the time limit is exceeded, VizAbility will display the message, “Answer: I’m sorry, but the process has been terminated because it took too long to arrive at an answer,” and will also suggest alternative questions to the user (see Section 4.5).
@@
D3
—To address contextual queries that require background or external information on what is available in the chart or its data, we have incorporated a Web Browser agent (ser, [n. d.]) to retrieve more general information relevant to chart comprehension. For example, when presented with the contextual query, “What do you mean by temperature anomalies,” the LLM responds with, “Temperature anomalies are any measure of temperatures that are unusual for a particular region, season, or time period. […]” Categorizing questions beforehand enabled us to streamline the process and eliminate unnecessary, resource-intensive prompts needed for analytical and visual queries.
@@
D4
—We seek to integrate users’ keyboard navigation with the conversational module via navigation queries. VizAbility currently supports two types of navigation queries: (a) wayfinding questions, in which, upon being provided a starting and ending point within the tree view, the model returns a series of directions dictating the shortest traversal and (b) orientation questions, in which the VizAbility returns the user’s current location within the tree view.
@@ 4.5.1. Proactive Error Mitigation Strategies
D6
—The versatility of VizAbility poses novel challenges in user query interpretation and answering. As reflected in the validation data set, some user queries may be too broad, ambiguous, or can even refer to variables that are not explicitly present in the data. For instance, the real user query “What kind of vaccine they used?” is irrelevant to the choropleth map displaying the share of the population that received at least one dose of the Covid-19 vaccine. User queries can also be poorly worded, such as “What parts of North America are not in the 80 to 100 percent range”, deeming it more difficult for an LLM to accurately interpret and compute the answer. These irrelevant and ambiguous queries can significantly decrease the overall accuracy of the system if not accounted for.
@@ 4.5.2. Reactive Error Mitigation Strategies
D6
—LangChain’s CSV Agent leverages chain of thought reasoning administered over a series of sequential prompts (in conjunction with pandas dataframe) to answer questions pertaining to a CSV file. Each individual prompt must therefore follow a specified format of Observation, Action, and Action Input for a subsequent prompt to properly parse the result (csv, [n. d.]). We have observed that sometimes an individual prompt in this sequential chain may be outputted in an incorrect format, thus triggering a cascading effect that ultimately results in an OutputParserException. To mitigate this error, we utilize prompt engineering and direct the CSV Agent Executor “You must follow the structure of Observation, Thought, Action, Action Input, etc.” to ensure that the output format is consistent between sequential queries. If the CSV Agent Executor is active for an extended period of time, we manually terminate it to preserve time-efficiency.
@@ 4.6. Fostering a More Accessible User Experience
D7
—Users can submit conversational queries via voice recordings that are processed via the Whisper speech recognition (whi, [n. d.]). However, oftentimes, enabling microphones can be problematic. Thus, we provide an alternative text box so that they can type the queries using the keyboard. Upon inputting their question (regardless of the modality), users are provided with an audible cue of “Loading. Please Wait”. Every subsequent 3 seconds, the user is exposed to yet another audible cue, this time “Still Loading”. This loading cue significantly improves transparency and mitigates any possible confusion that can arise from an unresponsive webpage.
@@
D7
—VizAbility does not solely display the answer, and instead provides the user query and brief justification behind its response in conjunction with the actual answer. For instance, the following is articulated by VizAbility when a user asks, “What is a choropleth map?”: “Your question ’What is a choropleth map?’ was categorized as being context-seeking, and as such, has been answered based on information found on the web.” By letting users know the scope of the answer (i.e., whether it was sourced from the internet, data, or the tree view), we allow users to verify and evaluate the effectiveness of the LLM response independently, thus bolstering user trust and system transparency.
@@
D7
—To help users figure out what queries are possible, VizAbility generates four initial queries, each of which belongs to one of the four query types. We achieve this suggestion by querying the LLM with a prompt that uses in-context impersonation (Salewski et al., 2023) (“Pretend you are a blind/low vision user who is presented with a chart.”), coupled with a text representation of the current chart. For instance, the LLM generates the following questions for the bar chart: “What is the temperature anomaly for the year 2020?”; “Can you provide a description of the color scheme used in the bar chart to represent the temperature anomalies?”; “Are there any patterns or relationships between the year and the temporal polarity of the temperature anomaly?”; “How do I get from my current position in the text representation to the x-axis?” These questions appear as interactive buttons, allowing users to either choose a suggested question or input their own.

### v4

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676414isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Visualization systems and tools
@@ 1. Introduction
Numerous assistive technologies have been developed to allow BLV users to access visualizations using sensory modalities other than vision (Kim et al., 2021). Tactile visualizations can provide a tangible representation of data while necessitating specialized hardware such as haptic displays  (Paneels and Roberts, 2009) and embossing machines (Engel and Weber, 2017a). On the other hand, sonification can enable users to discern trends and anomalies through sound (Wang et al., 2022), but it is typically limited to single-series data. Traditional methods for adapting web visualizations for screen readers, such as data tables and alternative text (Kim et al., 2021), often compromise the unique benefits of visualizations. Keyboard-based chart navigation (Weninger et al., 2015; Elavsky et al., 2024; Thompson et al., 2023; Zong et al., 2022) have emerged as an alternative solution. However, orienting oneself and navigating within complex chart encoding structures pose challenges for efficient data exploration (Kim et al., 2023a).
This work introduces VizAbility, a novel approach to augment keyboard navigation of chart content with conversational interaction (Figure 1). First, we use Olli’s tree view (referenced in (Blanco et al., 2022)) to create a keyboard-navigable text representation of a chart. Next, we enhance this tree view with an LLM-based question-and-answer module for addressing on-demand queries. These natural language queries allow users to understand the chart without needing to navigate and mentally synthesize different parts of the chart to derive insights. VizAbility uses the user’s position within the tree view to efficiently respond to both visual and analytical queries, facilitating the exploration of data trends and visual patterns. Moreover, VizAbility can handle contextual queries, providing background information about the chart, tailored specifically for the needs of BLV individuals (Kim et al., 2023b). We based these query types on actual questions asked by BLV people in previous studies (Kim et al., 2020, 2023b). Additionally, it can manage navigation queries, allowing users to control their position in the tree view through verbal commands.
@@
We carried out quantitative assessments to evaluate the question & answering pipeline. We evaluated response quality using a combined dataset of 979 real BLV user questions derived from previous research (Kim et al., 2023b) and 48 synthetically generated navigation queries. Splitting the dataset, 80% was used for testing and 20% for validation. Our query classification achieved an accuracy of 87.39%. The final response evaluation involved a manual assessment by two researchers, followed by a more scalable LLM-based evaluation using GPT4. Both the human and GPT4 assessments followed a 5-point Likert Scale that assigned a value ranging from “Very Poor” to “Very Good” depending on the response’s coherence to the ground truth. For the human evaluation, $77.36\%$ and $5.14\%$ of the responses were rated as “Very Good’ and “Good” respectively, contributing to an overall rate of $82.50\%$. We computed Kendall’s $\tau$ score to assess the consistency between the human and GPT4 assessment methods, observing significant alignment ($=0.5526$, $p<0.001$). As a baseline comparison, we used the GPT4-based evaluation to assess responses generated by GPT-4 with vision (GPT-4V) on the same test data. We observed the performance of GPT-4V was poorer, with only 27.96% and 10.10% of responses being designated as “Very Good” and “Good” respectively.
@@
  • Design and development of VizAbility, which incorporates an LLM-based question-and-answer module to enhance keyboard navigation of chart content for screen reader users.
@@
The VizAbility source code and dataset are available at https://dwr.bc.edu/vizability/.
@@ 2.1. Accessibility Systems for Data Visualization
Screen readers, utilizing text/speech modalities, stand as the predominant assistive technology, particularly for navigating web content. The go-to accessibility techniques for screen readers encompass alternative text and data tables. Yet, these strategies often reduce data visualizations to brief descriptions or mere numbers, undermining their inherent advantages. An alternative approach involves crafting keyboard-navigable text descriptions derived from the chart’s structure. A select group of data visualization tools and toolkits, such as HighCharts (Highcharts, 2022) and amCharts (amc, [n. d.]), offer some degree of this navigation and customization (Kim et al., 2023a). In recent times, several research efforts have advanced navigation capabilities, representing charts as traversable graph structures (Weninger et al., 2015; Godfrey et al., 2018; Zong et al., 2022; Thompson et al., 2023; Elavsky et al., 2024; Alam et al., 2023).
@@
To address the limitations of relying on a single sensory modality, multi-sensory perception is frequently utilized. A prevalent strategy involves merging verbal (speech) cues with non-verbal ones, such as sonification, tactile graphics, and haptic feedback. Examples include offering on-demand audio descriptions of touched elements (Goncu and Marriott, 2011; Gardner and Bulatov, [n. d.]; Landau and Gourgey, 2001) or pairing sonification with screen readers (Thompson et al., 2023; Taibbi et al., 2014) and braille (Seo et al., 2024). However, these solutions often necessitate specialized software and hardware, especially for interactive tactile support, making them expensive to implement.
@@ 2.2. Question & Answering Systems for Data Visualization
Recent studies have begun to focus on data visualization images (Hoque et al., 2022). For example, FigureQA (Kahou et al., 2018) offers a corpus tailored for yes/no questions, such as “Is Light Gold less than Periwinkle?”. Conversely, DVQA (Kafle et al., 2018) expands its purview to encompass questions about chart structure (“are the bars horizontal?”), data retrieval (“what percent of people prefer A?”), and reasoning (“Is A preferred more than B?”). While both FigureQA and DVQA rely on synthetically generated charts, PlotQA introduces a large-scale dataset of real-world scientific plots. Unlike the templated questions of the aforementioned datasets, ChartQA delivers human-composed questions, enhanced using LLMs (Masry et al., 2022). These models predominantly process pixel images as input. For instance, they extract data tables and other image features (Masry et al., 2022; Kantharaj et al., 2022), feeding them into vision and language task models (Cho et al., 2021). Consequently, their accuracy largely hinges on their image processing capabilities, often leading to suboptimal results (e.g., failing to recover data values due to the absence of data labels). In a different approach, Kim et al.(Kim et al., 2020) proposed a system that not only answers questions but also provides explanations, operating on Vega-lite(Satyanarayan et al., 2017) instead of images. All the current question-answering systems are limited to basic visualization types like bar, line, and pie charts.
@@ 3. VizAbility Design Decisions
Bridging the perceptual gap between BLV and sighted individuals requires a deep understanding of what the chart looks like. Previous research indicates that navigating charts based on their visual encoding helps BLV users gain a clearer visual understanding (Kim et al., 2023a). In this work, we use Olli (Blanco et al., 2022) to generate keyboard-navigable text representations of charts. This approach provides the visual structure of a chart in symbolic text form, helping answer visually oriented questions. Other keyboard-navigable text representations (e.g., a similar text structure extracted from HighCharts) would provide a similar benefit.
@@
D7: Follow standard accessibility principles
The W3C Web Accessibility Initiative specifies four essential principles for web accessibility: Perceivable, Operable, Understandable, and Robust (wai, 2023), while Chartability adds three more for visualization: Compromising, Assist-ive, and Flexible (Chartability Workbook, [n. d.]). Ensuring these principles in new assistive technology can be often overlooked. We strive to adhere to them in designing VizAbility. Examples include reaffirming user queries (P); suggesting cold-start queries (O); indicating delays in responses (U); allowing both speech and text inputs to formulate queries (R); accessing information via multiple modalities (C); converting data-centric language to user-friendly labels (A); and enabling switching between Olli and data table (F).
@@ 4.4. Rapid Chart Probing via Conversational Interaction
Keyboard navigation helps blind users grasp the chart’s visual structure and data, but can be cumbersome for higher-level comprehension, such as computing aggregates or understanding overall visuals. We integrate LLMs for intuitive chart question-answering and enhance interaction by combining keyboard navigation with speech inputs. Figure 3 shows the overview of the question-answering pipeline.
@@ 4.4.1. Data Set
The ground truths for the testing and validation sets were manually generated by two researchers independently and then merged by resolving conflicts. The process involved reading charts, calculating numbers, and searching for information online. Throughout the generation process, we emphasized verboseness. For instance, the ground truth response to the question “What is the vaccination rate of South Africa” is “The vaccination rate for South Africa is 36%”, as opposed to the more concise “36%”. The ground truth responses for navigation queries consist solely of starting and ending nodes to prioritize the model’s ability to translate ambiguous user wording into precise nodes within the tree view.
@@ 4.4.2. Supported Query Types
Analytical queriesD2 primarily focus on understanding the underlying data, such as “Is Africa the country that needs the vaccine the most?” or “What is the highest positive anomaly?” Visual queriesD2 relate to visual encoding information or demand visual interpretation, exemplified by questions like “What color is North America?” or “Is the line fluctuating?” Analytical and visual queries are not entirely distinct; visual queries often necessitate data interpretation, as in “Which country exhibits the darkest shades for both the lowest and highest values?”. Despite this overlap, we retain a conceptual separation in order to communicate to LLMs that a query involves information encoded visually and gain a more fine-grained understanding of VizAbility’s performance with chart-specific queries.
@@ 4.4.4. Query-Specific Prompting
D2Figure 6 illustrates our approach to handling analytical and visual queries. To circumvent the predefined token limit of the LLM, we consolidate the transformed data extracted from the Vega View (veg, [n. d.]) into an external CSV file. This file is then processed by LangChain’s CSV Agent (csv, [n. d.]), which operates in the background. Under the hood, this agent leverages the Pandas DataFrame agent, subsequently executing Python code generated by the LLM. We purposefully avoid including the entire raw dataset, recognizing that it might differ from the final view data. Often, the agent can get stuck in an infinite loop of thinking. To prevent this issue, we have implemented a time constraint. If the time limit is exceeded, VizAbility will display the message, “Answer: I’m sorry, but the process has been terminated because it took too long to arrive at an answer,” and will also suggest alternative questions to the user (see Section 4.5).
@@
D3—To address contextual queries that require background or external information on what is available in the chart or its data, we have incorporated a Web Browser agent (ser, [n. d.]) to retrieve more general information relevant to chart comprehension. For example, when presented with the contextual query, “What do you mean by temperature anomalies,” the LLM responds with, “Temperature anomalies are any measure of temperatures that are unusual for a particular region, season, or time period. […]” Categorizing questions beforehand enabled us to streamline the process and eliminate unnecessary, resource-intensive prompts needed for analytical and visual queries.
@@
D4—We seek to integrate users’ keyboard navigation with the conversational module via navigation queries. VizAbility currently supports two types of navigation queries: (a) wayfinding questions, in which, upon being provided a starting and ending point within the tree view, the model returns a series of directions dictating the shortest traversal and (b) orientation questions, in which the VizAbility returns the user’s current location within the tree view.
@@ 4.5.1. Proactive Error Mitigation Strategies
D6—The versatility of VizAbility poses novel challenges in user query interpretation and answering. As reflected in the validation data set, some user queries may be too broad, ambiguous, or can even refer to variables that are not explicitly present in the data. For instance, the real user query “What kind of vaccine they used?” is irrelevant to the choropleth map displaying the share of the population that received at least one dose of the Covid-19 vaccine. User queries can also be poorly worded, such as “What parts of North America are not in the 80 to 100 percent range”, deeming it more difficult for an LLM to accurately interpret and compute the answer. These irrelevant and ambiguous queries can significantly decrease the overall accuracy of the system if not accounted for.
@@ 4.5.2. Reactive Error Mitigation Strategies
D6—LangChain’s CSV Agent leverages chain of thought reasoning administered over a series of sequential prompts (in conjunction with pandas dataframe) to answer questions pertaining to a CSV file. Each individual prompt must therefore follow a specified format of Observation, Action, and Action Input for a subsequent prompt to properly parse the result (csv, [n. d.]). We have observed that sometimes an individual prompt in this sequential chain may be outputted in an incorrect format, thus triggering a cascading effect that ultimately results in an OutputParserException. To mitigate this error, we utilize prompt engineering and direct the CSV Agent Executor “You must follow the structure of Observation, Thought, Action, Action Input, etc.” to ensure that the output format is consistent between sequential queries. If the CSV Agent Executor is active for an extended period of time, we manually terminate it to preserve time-efficiency.
@@ 4.6. Fostering a More Accessible User Experience
D7—Users can submit conversational queries via voice recordings that are processed via the Whisper speech recognition (whi, [n. d.]). However, oftentimes, enabling microphones can be problematic. Thus, we provide an alternative text box so that they can type the queries using the keyboard. Upon inputting their question (regardless of the modality), users are provided with an audible cue of “Loading. Please Wait”. Every subsequent 3 seconds, the user is exposed to yet another audible cue, this time “Still Loading”. This loading cue significantly improves transparency and mitigates any possible confusion that can arise from an unresponsive webpage.
@@
D7—VizAbility does not solely display the answer, and instead provides the user query and brief justification behind its response in conjunction with the actual answer. For instance, the following is articulated by VizAbility when a user asks, “What is a choropleth map?”: “Your question ’What is a choropleth map?’ was categorized as being context-seeking, and as such, has been answered based on information found on the web.” By letting users know the scope of the answer (i.e., whether it was sourced from the internet, data, or the tree view), we allow users to verify and evaluate the effectiveness of the LLM response independently, thus bolstering user trust and system transparency.
@@
D7—To help users figure out what queries are possible, VizAbility generates four initial queries, each of which belongs to one of the four query types. We achieve this suggestion by querying the LLM with a prompt that uses in-context impersonation (Salewski et al., 2023) (“Pretend you are a blind/low vision user who is presented with a chart.”), coupled with a text representation of the current chart. For instance, the LLM generates the following questions for the bar chart: “What is the temperature anomaly for the year 2020?”; “Can you provide a description of the color scheme used in the bar chart to represent the temperature anomalies?”; “Are there any patterns or relationships between the year and the temporal polarity of the temperature anomaly?”; “How do I get from my current position in the text representation to the x-axis?” These questions appear as interactive buttons, allowing users to either choose a suggested question or input their own.